OpenStudy (anonymous):

What is the simplified form of x-3/x^2+x-12 •x+4/x^2+8x+16 ? Help!

OpenStudy (anonymous):

what is that dot for?

OpenStudy (anonymous):

factor the denominators so that the numerator will cancel

OpenStudy (anonymous):

multiplication

OpenStudy (anonymous):

the final answer is \[\left( x + 4 \right)^{-2} = \frac{ 1 }{ \left( x + 4 \right)^{2} }\]

OpenStudy (anonymous):